# Certificate-Sensitive Subset Sum: Realizing Instance Complexity

### v4

@@ 5.1 Double Meet-in-the-Middle Solver
The runtime follows by applying the analysis of Theorem 1 to each of the two halves of size $n/2$. The total work is governed by the larger of the two resulting enumerations due to the solver’s interleaved execution model, as justified in the subsequent section. ∎
5.2 Justification for the $\max(U_{0},U_{1})$ Runtime Bound in the Interleaved Model
Classical meet-in-the-middle solvers construct full enumeration tables for both halves of the input, yielding a runtime proportional to $U_{0}+U_{1}$. In contrast, our solver executes an interleaved enumeration and merge process, where elements of $\Sigma(\ell_{0})$ and $\Sigma(\ell_{1})$ are matched on-the-fly. We now justify the tighter runtime bound of $O(\max(U_{0},U_{1})\cdot n^{2})$ for the deterministic variant and $O(\max(U_{0},U_{1})\cdot n)$ for the randomized one.
Naive Bound.
A sequential evaluation that first constructs both $\Sigma(\ell_{0})$ and $\Sigma(\ell_{1})$ in full would incur a cost of $O((U_{0}+U_{1})\cdot n^{2})$ in the deterministic model.
Stronger Claim.
Our solver avoids full tabulation and instead performs feasibility checks incrementally. The merge logic operates online, emitting each candidate from one half and immediately testing it against the current frontier from the other. This yields a tighter complexity bound:
$T(n)=O(\max(U_{0},U_{1})\cdot n^{2})$
in the deterministic model, and
$T(n)=O(\max(U_{0},U_{1})\cdot n)$
in the randomized setting, where canonicality checks are $O(1)$ in expectation.
Formal Justification.
The total computational cost arises from:
  • enumerating all $U_{i}$ unique subset sums in side $\ell_{i}$, with $O(n)$ cost per bitmask update,
  • executing a constant number of hash table or lookup operations per candidate sum,
  • performing canonical deduplication when necessary.
Because the algorithm proceeds in rounds, extending one side’s frontier and probing the other immediately, the overall control flow is governed by the larger of $U_{0}$ or $U_{1}$. The smaller half’s contribution is subsumed within the dominant half’s traversal. Thus, no full duplication of effort occurs, and the interleaving guarantees that the additive bound $U_{0}+U_{1}$ is overly pessimistic.
Contrast with Horowitz–Sahni.
Classical meet-in-the-middle (e.g., Horowitz–Sahni [8]) uses an offline, two-phase architecture: first build both lists, then merge. This necessitates $O(U_{0}+U_{1})$ full enumeration regardless of structure. Our model, by contrast, halts exploration early when feasible and avoids redundant enumeration, exploiting the fact that $\Sigma(\ell_{0})$ and $\Sigma(\ell_{1})$ need not be fully materialized to solve the instance.
Conclusion.
The interleaved solver’s runtime is structurally governed by the maximum work performed on either half. This aligns with both our theoretical analysis and empirical benchmarks, and further emphasizes the adaptivity of IC-SubsetSum to asymmetric structure.
5.3 Worst-Case Runtime via Controlled Aliasing
In addition to its adaptive performance, IC-SubsetSum guarantees a strict worst-case improvement over Horowitz–Sahni, even on collision-free instances. This is achieved via a deterministic Controlled Aliasing rule: a structural redundancy technique that reduces the enumeration space without sacrificing correctness. We describe an enhanced version of this technique that ensures full correctness for both the decision and construction variants of Subset–Sum.
Aliasing Rule.
Let $S$ be split into halves $\ell_{0}$ and $\ell_{1}$. In each half, select a pair of distinct elements—an alias pair—e.g., $(x_{0},x_{1})\in\ell_{0}$. During enumeration, we syntactically substitute the value of $x_{1}$ with that of $x_{0}$:
$\text{If }x_{1}\in P,\quad\text{compute }\sigma(P)\text{ using }x_{0}\text{ in place of }x_{1}.$
This substitution affects only the computed value; the subset’s bitmask identity is preserved. The result is that multiple distinct subsets collapse onto the same sum, reducing the enumeration state space. As proven in Appendix E, this yields a bound of
$O^{*}\left(\tfrac{3}{4}\cdot 2^{n/2}\right)=O^{*}(2^{n/2-\log_{2}(4/3)})\approx O^{*}(2^{n/2-0.415}).$
Canonical Aliasing via Parallel Memoization Lanes.
To preserve correctness under aliasing, we introduce a 4-lane memoization model. Each sum $\sigma$ is now associated with four parallel canonical entries, one for each aliasing state of the subset that generated it:
  1. (i)
    Neither $x_{0}$ nor $x_{1}$ included,
  2. (ii)
    $x_{0}$ included only,
  3. (iii)
    $x_{1}$ included only,
  4. (iv)
    Both $x_{0}$ and $x_{1}$ included.
These lanes are disjoint: a new prefix competes only within its own lane for lexicographic minimality. Thus, all semantically distinct witnesses are retained. The memoization structure becomes:
$\texttt{Memo}[\sigma]\in\left(\{\bot\}\cup\{0,1\}^{n}\right)^{4},$
i.e., a 4-tuple of bitmasks per reachable sum.
Compensatory Expansion.
During the final merge phase, we evaluate all valid interpretations of aliased sums. Let $s_{L}\in\Sigma(\ell_{0})$ and $s_{R}\in\Sigma(\ell_{1})$ be enumerated sums, and let $(x_{0},x_{1})$ and $(y_{0},y_{1})$ be the alias pairs in $\ell_{0}$ and $\ell_{1}$, respectively.
Define corrected sums:
$s^{\prime}_{L}$ $=s_{L}-x_{0}+x_{1}\quad\text{(if $x_{1}$ was present)}$)
$s^{\prime}_{R}$ $=s_{R}-y_{0}+y_{1}\quad\text{(if $y_{1}$ was present)}.$) .
We check all combinations:
$s_{L}+s_{R}=t,\quad s^{\prime}_{L}+s_{R}=t,\quad s_{L}+s^{\prime}_{R}=t,\quad s^{\prime}_{L}+s^{\prime}_{R}=t.$
Witness Reconstruction.
Each of the four lanes retains a canonical prefix corresponding to its aliasing state, allowing for unambiguous witness recovery by combining the correct bitmasks from each half.
Remark 2.
While we focus on aliasing pairs for simplicity, the underlying 4-lane correctness mechanism is a scalable principle. The technique can be extended by using multiple aliases to the same element (e.g., $x_{1}\to x_{0}$, $x_{2}\to x_{0}$), which can yield a greater search space reduction. For example, a trio of aliased elements provides a speedup of $\epsilon=1$. Each additional alias increases the number of “lanes” and the complexity of the compensatory checks by a constant factor, but the asymptotic complexity remains unchanged. Analyzing the optimal trade-off is a promising direction for future work.
6 Refined View of Fine-Grained Hardness
Our algorithm challenges the standard landscape of fine-grained reductions. It shows that many hardness assumptions implicitly rely on a promise of low collision entropy—a structural assumption that classical reductions neither justify nor quantify. We make this dependence precise.
6.1 The Collision-Free Promise
We begin by formalising a restricted slice of Subset–Sum where classical hardness results are still meaningful.
Definition 3 (Collision-Resistant Subset–Sum).
For a parameter $\delta\in(0,1]$, define the class of instances with high entropy as
@@ 6.1 The Collision-Free Promise
This condition ensures that any adaptive or certificate-sensitive algorithm cannot short-circuit the instance via early success or structural pruning. In particular, it guarantees that no-solution instances still require a near-exhaustive traversal of the exponentially large certificate space. Thus, hardness is preserved not just in satisfiable cases, but in verifying infeasibility—a setting where adaptive enumeration offers no shortcut unless entropy is low.
6.2 Refining SETH Lower Bounds
@@
Theorem 5 (Revised SETH Lower Bound under Entropy Constraints).

### v6

@@ 6.2 Refining SETH Lower Bounds
Take a collision-resistant reduction from $k$-SAT that maps instances to $(S,t)\in\textsc{Subset--Sum}^{\geq\delta}$ with polynomial blowup. If an $O^{*}(2^{\epsilon n})$ algorithm existed for all such $(S,t)$ with $\epsilon<\delta$, composing would yield an $O^{*}(2^{\epsilon^{\prime}m})$ algorithm for $k$-SAT with $\epsilon^{\prime}<1$, contradicting SETH. ∎
7 Proof-of-Concept Experiments
We provide proof-of-concept experiments to sanity-check the theory. The goal is not a systems study but to verify that the observed running time tracks the structural parameter that our analysis identifies. We benchmark on synthetic instances where we can precisely control structure, using $n=48$ so that full per-half enumeration is feasible on a commodity desktop CPU. Throughout we split $S$ into two halves by alternating indices, so each half has size $k=n/2=24$. For each instance we fully enumerate both halves to obtain $U_{0}:=|\Sigma(\ell_{0})|$ and $U_{1}:=|\Sigma(\ell_{1})|$, and we measure wall-clock time of the deterministic enumerator (hashing disabled) to avoid randomness. Our hypothesis is that introducing additive redundancy collapses $U_{i}$, and runtime scales proportionally with $U_{0}+U_{1}$ (hence with $\widehat{U}:=\max\{U_{0},U_{1}\}$ for fixed $n$).
We manipulate three knobs known to induce subset-sum collisions: numeric density, duplicate elements, and additive progressions. For each setting we run the full certificate construction on each half (no early termination) to measure the cost of generating $\Sigma(\ell_{i})$. Additional methodology appear in Appendix D.
  • Numeric density. Restricting values to a smaller bit-length $w$ forces collisions by the pigeonhole principle.
  • Duplicate elements. Introducing identical elements creates trivial additive dependencies.
  • Additive progressions. Planting short arithmetic progressions creates correlated dependencies.
Table 1: Effect of structure on the per-half distinct-sum count. We report the ratio $U_{i}/2^{k}$ with $k=n/2=24$ (median over halves for a representative run).
Structural knob Setting $U_{i}/2^{k}$
Numeric density (n=48n{=}48) $w=32\to 24\to 16$ $1.00\to 0.84\to 0.027$
Duplicates (n=48n{=}48) $0\to 2\to 4$ duplicates $1.00\to 0.57\to 0.32$
Additive progressions (n=48n{=}48) 1 seq (len 4) / 2 seq (len 4) $0.69$ / $0.47$
Results (sanity check).
We observe large, systematic variation in $U_{i}$ under these perturbations: uniform-like inputs typically satisfy $U_{i}\approx 2^{k}$, whereas structured inputs collapse by orders of magnitude. For fixed $n$, the measured runtime of IC-SubsetSum varies proportionally with $U_{0}+U_{1}$ (equivalently with $\widehat{U}$ up to a factor 2), consistent with the proven $O\!\big{(}(U_{0}{+}U_{1})\,n^{2}\big{)}$ bound. These experiments corroborate the central claim that structure (collisions) $\Rightarrow$ smaller $U$ $\Rightarrow$ faster enumeration.
8 Related Work
Subset Sum Algorithms.
The classical Horowitz–Sahni meet-in-the-middle algorithm [8] remains the standard worst-case baseline for Subset–Sum, with $O^{*}(2^{n/2})$ runtime. A space-time refinement due to Schroeppel and Shamir achieves the same time with $O^{*}(2^{n/4})$ space [14]. Despite extensive efforts, no deterministic unconditional improvement below $2^{n/2}$ was known in general [4, 17]. Our approach departs from this tradition in two ways: (i) we analyze runtime as a function of the number of distinct subset sums $U$ rather than just $n$, yielding certificate-sensitive bounds; and (ii) via Controlled Aliasing under CNF with compensatory merge, we obtain a deterministic worst-case improvement.
A recent randomized algorithm by Bringmann, Fischer, and Nakos [4] gives the first unconditional improvement over Bellman’s dynamic programming for the target-constrained decision problem, running in $\widetilde{O}(|S(X,t)|\cdot\sqrt{n})$ time. Their technique focuses on reachable sums up to a target $t$ via algebraic/combinatorial methods. By contrast, our algorithm deterministically enumerates the full set of unique subset sums $\Sigma(S)$, supports canonical constructive witnesses, and adapts to instance structure. The two lines are complementary: they address different problem variants and performance measures. We also note pseudo-polynomial improvements for dense/structured instances (e.g., [9, 5]) that are orthogonal to our certificate-size viewpoint.
Instance Complexity and Adaptive Algorithms.
Instance Complexity was introduced by Orponen–Ko–Schöning–Watanabe [10] to capture the inherent difficulty of individual instances. Although potent as a conceptual tool, IC was long viewed as non-constructive (see, e.g., [6]). Our results give a constructive realization: we generate the IC certificate online with runtime scaling in the certificate size $U$, and we wrap it to satisfy the partial-correctness conventions of IC (see §4.3).
Fine-Grained Complexity.
The fine-grained framework [16] provides tight conditional barriers (e.g., under SETH). We add a structural dimension: such lower bounds for Subset–Sum implicitly target collision-resistant instances with near-maximal $U$. Making this dependency explicit complements work on compressibility-aware reductions [1]; our formulation via $\Sigma(S)$ and collision entropy is instance-level and operational.
Collision Structure in Combinatorics.
The role of collisions in subset sums appears in additive combinatorics and cryptanalysis. Austrin-Kaski-Koivisto-Nederlof [2] analyze structure under anti-concentration, while Howgrave-Graham and Joux [7] exploit collisions via randomized meet-in-the-middle techniques to obtain sub-$2^{n/2}$ behavior in certain regimes. Classic Littlewood–Offord–type results [15] bound collision counts via anti-concentration. Our contribution is algorithmic and instance-sensitive: we prune duplicate sums in real time and link entropy collapse directly to runtime; and, distinct from prior randomized collision tricks, our Controlled Aliasing (under CNF with compensatory merge) is a deterministic, target-independent transformation that preserves correctness while provably shrinking the explored state space.
9 Future Directions
Closing Algorithmic Gaps for Subset Sum.
Two natural questions remain. First, for the randomized $O(U\cdot n)$ algorithm, can the linear factor in $n$ be removed to obtain $\tilde{O}(U)$ time? Second, and more fundamentally, can the deterministic $O(U\cdot n^{2})$ bound be improved to match the randomized $O(U\cdot n)$, e.g., via a deterministic canonicality test that avoids the $O(n)$ lexicographic comparison cost while preserving the local model in Conjecture 1?
Improving the Controlled Aliasing Bound.
Our worst-case speedup $\varepsilon=\log_{2}(\frac{4}{3})$ arises from a single alias pair per half under the 2-lane canonicalization keyed by $(\tilde{\sigma},\chi)$ (CNF). Straightforward pattern-based generalizations for alias groups of size $g$ induce $2^{g}$ lanes and exponential overhead, keeping $\varepsilon$ constant. We conjecture that a count-based scheme—tracking only the number of chosen elements from each alias group (yielding (g+1)(g{+}1) buckets) with polynomial overhead—could permit $g=\Theta(\log n)$, improving the multiplicative reduction from $2^{g}$ to (g+1)(g{+}1). In principle, this could yield $\varepsilon=\Theta(\log n)$, i.e., a worst-case runtime of $2^{n/2}/n^{c}$ for some constant $c>0$. Making this rigorous requires (i) a safety proof for count-only canonicalization (lane invariant and merge completeness), and (ii) a certified reconstruction path from counts to witnesses without reintroducing exponential blowup.
Structure-Sensitive Runtimes in NP.
Which other NP-complete problems admit certificate-sensitive runtimes? Collision entropy provides a structural knob complementary to solution density. For Partition, Knapsack, and Bin Packing, one may aim to construct compressed summaries of feasible configurations whose construction time is proportional to summary size. Formalizing such certificates and proving instance-sensitive enumeration bounds are promising directions.
Derandomization of Local Canonicality.
The randomized $O(U\cdot n)$ bound uses 2-universal hashing to filter collisions so that lexicographic comparisons are needed only rarely (expected $O(1)$ per check). Is there a deterministic analogue achieving sublinear-time canonicality tests in the local deduplicating model? Such a result would match the conjectured optimal runtime (Conjecture 1) and yield a practical local derandomization.
Constructive Instance Complexity Beyond Subset Sum.
We showed that the certificate $\Sigma(S)$ can be constructed in deterministic $O(U\cdot n^{2})$ time (or expected $O(U\cdot n)$). Can analogous online certifiers be built for problems such as 3SAT or Clique under appropriate entropy/compressibility assumptions? A broader theory of algorithmic IC—tracking both information content and construction cost—remains to be developed.
Toward a Complexity Class for Certificate-Sensitive Problems.
Our solver runs in time polynomial in the input size $n$ and in the size of an intrinsic instance-specific certificate ($U=|\Sigma(S)|$). This suggests a class we tentatively call P-IC (Polynomial in Instance Certificate): problems for which a natural instance certificate can be constructed and then used to solve the instance in time polynomial in the certificate size and input size. Pinning down robust certificate notions (e.g., sets of satisfying assignments for SAT, achievable value/weight pairs for Knapsack) and charting the boundaries of P-IC is an intriguing direction for future work.
10 Conclusion
Theoretical Advance.
This paper develops a framework for structure-aware enumeration in NP-complete problems, using Subset Sum as a case study. We introduce the certificate-sensitive parameter $U:=|\Sigma(S)|$ and show that it aligns with both per-instance difficulty and observed runtime. The results help explain why many real-world instances are easier, why SETH-based reductions can be structurally brittle, and how instance complexity can be made constructive.
Algorithmic Advance.
IC-SubsetSum delivers a set of guarantees that advance the state of the art:
    Certificate-sensitive runtime.
  • Certificate-sensitive runtime. A deterministic worst-case bound of $O(U\cdot n^{2})$ and an expected randomized bound of $O(U\cdot n)$, realizing a constructive link to Instance Complexity by tying performance to the certificate size $U$, which is discovered online in an anytime fashion.
  • Worst-case improvement. Via Controlled Aliasing under count-based canonical normalization (CNF) at the alias indices, a deterministic bound of $O^{*}(2^{\,n/2-\varepsilon})$ with $\varepsilon=\log_{2}(\frac{4}{3})\approx 0.415$, improving on the classical $O^{*}(2^{n/2})$ Horowitz–Sahni bound for all sufficiently large $n$.
  • Real-time certificate generation. Online construction of the constructive certificate, storing the lex-minimal witness for each $\sigma\in\Sigma(S)$ without duplication.
Closing the gap between the deterministic and randomized bounds remains a key open challenge, as does the question of whether an $\tilde{O}(U)$ runtime is achievable. [^1] May all your programs be short and quick (Fortnow [6]) and may their runtimes reveal the structure within.
Appendix Overview
This appendix expands on the enumeration logic, correctness conditions, and structural optimizations introduced in Sections 46. It is organized as follows:
  • Appendix A details the column-wise subset-sum enumerator, including the suppression-on-overwrite invariant and the lex-minimal canonicalization rule, with full pseudocode.
  • Appendix B presents the interleaved double–meet-in-the-middle framework, the real-time Check/CheckAliased procedures (including the $2\times 2$ compensatory merge under aliasing), and the work accounting used in the certificate-sensitive bounds.
  • Appendix C states the algebraic combination lemmas used by the solver (direct, complement, and mixed cases) and the associated correctness checks for sum recombination that power the real-time Check routines.
  • Appendix D provides additional experimental data, including plots of runtime scaling versus $U$ and measured collision entropy.
  • Appendix E gives the full Controlled Aliasing framework: the 2-lane $(\tilde{\sigma},\chi)$ memoization with affine correction, the count-based canonical normalization (CNF) at the alias indices, and the expansion policy. It includes the key lemmas—Lemma 8, Lemma 6, Lemma 7, Lemma 9, Lemma 10—and the proof of Theorem 6.
Appendix A Column-wise Enumerator: Full Details
Algorithm Description
Column-wise expansion.
We organize the search by columns of increasing subset size. The state space consists of canonical partial sums $\sigma$ paired with their current prefix (bitmask) $R$. At column $k$ we extend each canonical state of size k1k{-}1 by adding one unused element $a_{i}\notin R$, producing $\sigma^{\prime}=\sigma+a_{i}$ with prefix $R^{\prime}=R\cup\{a_{i}\}$. If $\sigma^{\prime}$ is new, or if $R^{\prime}$ is lexicographically smaller than the current representative, we update the representative of $\sigma^{\prime}$. This yields at most $O(U\cdot n)$ total extension attempts from canonical states. As established in Section 4.1, the overall complexity is governed by the cost of canonicality checks during collisions.
High-density handling.
To guarantee correctness in high-collision regimes, we impose a deterministic canonicalization based on lexicographic order of index bitmasks: scan indices $1\!\to\!n$ and prefer $0<1$ at the first differing position. Each numeric sum $\sigma$ keeps exactly one canonical representative (lex-minimal mask), which prevents non-deterministic pruning and ensures a single expansion path per sum.
4-Column Litmus Test.
The real-time collision checks enable a quick hardness probe. By running only the first $k$ columns (e.g., $k=4$) in polynomial time $O(n^{k})$, we empirically estimate the early collision rate. High early collision typically predicts small $U$, signaling an “easy” instance; low early collision predicts a large $U$, signaling a harder, unstructured input. This offers a lightweight pre-check before any full exponential exploration.
Reader’s Guide (what to look for).
  • Separation of concerns: (i) enumeration of prefixes, (ii) canonical tie-breaking, (iii) suppression-on-overwrite. Any scheduler (BFS/DFS/priority) that respects these yields the same canonical set.
  • Spec vs. implementation: We first give a minimal recursive specification that makes the logic transparent; the column-wise iterative code that follows is the optimized implementation used in our analysis.
  • Where the $O(U\!\cdot\!n^{2})$ comes from: $O(U\!\cdot\!n)$ extension attempts times $O(n)$ lex-compare cost on collisions, with suppression ensuring each canonical state expands at most once.
Minimal Recursive Specification (for clarity)
Algorithm 2 EnumerateUniqueSubsetSums-Recursive (spec; with suppression-on-overwrite)
[…]
2:Memo maps each numeric sum to its canonical (lex-min) bitmask
3:Memo { $\leftarrow\{\,0\mapsto\texttt{empty\_bitmask}\,\}$; doNotExtend[empty_bitmask] $\leftarrow$ false
4:procedure Recurse($\texttt{mask},\,\texttt{sum},\,\texttt{next}$) $\triangleright$ include-only-forward indices ensure a unique generation path
5:  if doNotExtend[maskthen return
6:  for $i=\texttt{next}$ to $n$ do
7:   $\texttt{new\_sum}\leftarrow\texttt{sum}+a_{i}$; $\texttt{new\_mask}\leftarrow\textsc{SetBit}(\texttt{mask},i)$
8:   if $\texttt{new\_sum}\notin\texttt{Memo}$ then
9:     Memo[new_sum] $\leftarrow\texttt{new\_mask}$; doNotExtend[new_mask] $\leftarrow$ false
10:     Recurse($\texttt{new\_mask},\,\texttt{new\_sum},\,i{+}1$)
11:   else if IsLexicographicallySmaller(new_mask, Memo[new_sum]then
12:     $\texttt{old}\leftarrow\texttt{Memo}[\texttt{new\_sum}]$; doNotExtend[old] $\leftarrow$ true $\triangleright$ suppress-on-overwrite
13:     Memo[new_sum] $\leftarrow\texttt{new\_mask}$; doNotExtend[new_mask] $\leftarrow$ false
14:     Recurse($\texttt{new\_mask},\,\texttt{new\_sum},\,i{+}1$)
15:Recurse($\texttt{empty\_bitmask},\,0,\,1$); return Memo
Mechanics of Frontier/Rep/doNotExtend (matches Inv. 1)
  • Frontier holds exactly the canonical states scheduled to be expanded next. We advance by columns (subset size), so each state in Frontier is expanded at most once per Inv. 1.
  • Rep[ $\sigma$ ] points to the current representative state for numeric sum $\sigma$ in this column’s scheduler. It lets us (i) find and (ii) immediately suppress a displaced representative when a lex-smaller mask appears.
  • doNotExtend is a per-state flag. When a representative is overwritten by a lex-smaller mask, we set doNotExtend on the displaced state. Even if a displaced state were still present in some queue, the flag enforces no further expansion, ensuring each canonical sum expands exactly once.
  • Optional unscheduling. We also remove the displaced state from the current NextFrontier if it was tentatively enqueued, which avoids a wasted dequeue in this same column. This is a micro-optimization; correctness relies only on doNotExtend.
A.1 Iterative Column-wise Implementation (optimized)
Algorithm 3 EnumerateUniqueSubsetSums (with suppression-on-overwrite)
1:Set $S=\{a_{1},\dots,a_{n}\}$ (indices $1..n$)
2:Memo maps each numeric sum to its canonical (lex-min) bitmask
@@ A.1 Iterative Column-wise Implementation (optimized)
6:for $k=1$ to $n$ do $\triangleright$ BFS by columns (subset size); each canonical state expands at most once
7:  $\texttt{NextFrontier}\leftarrow\emptyset$
8:  for all $\texttt{state}\in\texttt{Frontier}$ do
9:   if state.doNotExtend then continue    $\triangleright$ Inv. 1: suppressed reps never expand
10:   for $i=1$ to $n$ do
11:     if bit $i$ is not set in state.mask then
12:      $\texttt{new\_sum}\leftarrow\texttt{state.sum}+a_{i}$
13:      $\texttt{new\_mask}\leftarrow\textsc{SetBit}(\texttt{state.mask},i)$
14:      if $\texttt{new\_sum}\notin\texttt{Memo}$ then $\triangleright$ first time we see this numeric sum: install canonical rep and schedule it
15:        $\texttt{Memo[new\_sum]}\leftarrow\texttt{new\_mask}$
16:        $\texttt{new\_state}\leftarrow\texttt{State}(\texttt{new\_sum},\texttt{new\_mask},\texttt{false})$
17:        $\texttt{Rep[new\_sum]}\leftarrow\texttt{new\_state}$
18:        $\texttt{NextFrontier}\leftarrow\texttt{NextFrontier}\cup\{\texttt{new\_state}\}$
19:      else if IsLexicographicallySmaller(new_mask, Memo[new_sum]then $\triangleright$ lex-smaller: overwrite the rep and suppress the displaced one (Inv. 1)
20:        if $\texttt{Rep[new\_sum]}\neq\texttt{nil}$ then
21:         $\texttt{Rep[new\_sum].doNotExtend}\leftarrow\texttt{true}$ $\triangleright$ suppress old representative from any future expansion
22:         /* optional micro-optimization: if scheduled, unschedule it in this column */
23:         remove Rep[new_sum] from NextFrontier if present
24:        $\texttt{Memo[new\_sum]}\leftarrow\texttt{new\_mask}$ $\triangleright$ install the new canonical representative
25:        $\texttt{new\_state}\leftarrow\texttt{State}(\texttt{new\_sum},\texttt{new\_mask},\texttt{false})$
26:        $\texttt{Rep[new\_sum]}\leftarrow\texttt{new\_state}$ $\triangleright$ Rep now points at the (unique) expandable canonical state
27:        $\texttt{NextFrontier}\leftarrow\texttt{NextFrontier}\cup\{\texttt{new\_state}\}$
@@
29:  $\texttt{Frontier}\leftarrow\texttt{NextFrontier}$ $\triangleright$ advance one column; each canonical sum’s representative is expanded at most once
30:  if $\texttt{Frontier}=\emptyset$ then break
31:return Memo
Algorithm 4 IsLexicographicallySmaller (1-indexed, $0<1$)
1:function IsLexicographicallySmaller(AA, $B$) $\triangleright$ AA, $B$ are $n$-bit masks
2:  for $i=1$ to $n$ do
3:   $a_{i}\leftarrow$ bit $i$ of $A$; $b_{i}\leftarrow$ bit $i$ of $B$
4:   if $a_{i}\neq b_{i}$ then return $(a_{i}<b_{i})$    $\triangleright$ prefer 0 at first difference
5:  return false $\triangleright$ equal masks
Algorithm 5 SetBit (helper)
1:function SetBit($M,i$)
2:  return $M$ with bit $i$ set to 11
A.2 Implementation Invariants
To guarantee correctness and performance, the enumerator maintains:
  • Deterministic tie-breaking. For any numeric sum $\sigma$, the lex-minimal bitmask (Algorithm 4) is the unique representative. This ensures global canonicity across columns.
  • Efficient storage. We store at most $U$ representatives, each as an $n$-bit mask, for total space $O(U\cdot n)$.
  • Duplication elimination. A numeric sum may be discovered multiple times, but its canonical representative is expanded at most once. All non-canonical discoveries either lose the tie or overwrite and suppress the previous representative.